Julie, Bill, and Travis have agreed to distribute the profits of JBT Partners in the ratio of 3:2:5. If the firm made a net profit of $81,000 this year, how much of this $81,000 will Travis receive?
a. $27,000.00
b. $24,300.00
c. $8,100.00
d. $40,500.00


Answer by checkley71(8403)

SEEING AS THERE ARE 2=3+5=10 SHARES & TRAVIS HAS 5 OF THE 10 THEN TRAVIS GETS 5/10 OR 50% OF THE PROFIT
50% OT 81000 IS .5*81000=$40,500 ANSWER d)